# SiamSeg: Self-Training with Contrastive Learning for Unsupervised Domain Adaptation Semantic Segmentation in Remote Sensing

### v2

@@ III-D Contrastive Learning
$L_{CLR}^{i}=\frac{1}{2}\cdot D(p_{1}^{(i)},stop\_grad(z_{2}^{(i)}))+\frac{1}{2}\cdot D(p_{2}^{(i)},stop\_grad(z_{1}^{(i)})).$ (7)
In the loss function, the stop-gradient operation is a crucial step, treating the variables in $stop\_grad(\cdot)$ as constants. This operation effectively prevents representation collapse [44]. Specifically, in the first term of Equation (7), the encoder processing $x_{T_{2}}^{(i)}$ does not receive gradients from $z_{2}^{(i)}$, while in the second term, the encoder processing $x_{T_{1}}^{(i)}$ does receive gradients from $p_{2}^{(i)}$. This design effectively enhances the model’s representation learning capabilities, improving its performance in the target domain.
III-E Proposed UDA Losses
In a given training step, we acquire the $m$-th image and its corresponding label from the source domain, represented as $x_{S}^{(m)}$ and $y_{S}^{m}$. Simultaneously, we obtain the $m$-th image from the target domain and its pseudo-label generated by the teacher network $t_{\theta}$, denoted as $x_{T}^{(n)}$ and $p_{T}^{(n)}$. The total loss function we mathematically define is as follows:
$L_{total}^{m;n}=L_{S}^{(m)}+\beta\cdot L_{T}^{(n)}+\gamma\cdot L_{CLR}^{(n)}.$ (8)
@@ III-E Proposed UDA Losses
Specifically, the loss $L_{S}^{(m)}$ guides the model’s performance on source domain data, while $L_{T}^{(n)}$ leverages pseudo-labels to drive learning in the target domain, further enhancing the model’s generalization ability in this domain. Finally, $L_{CLR}^{(n)}$ introduces richer feature representations through contrastive learning, enabling the model to better adapt to potential differences between the source and target domains. By integrating these losses, we can effectively reduce the domain gap between the source and target domains, thereby improving the model’s performance on the unlabeled target domain data.
III-F SiamSeg Network Architecture
III-F1 overall
Figure 2: The SiamSeg network $g_{\theta}$ is designed for image segmentation and comprises a feature extraction backbone $f$ and a decoding head. An Exponential Moving Average (EMA) teacher network $t_{\theta}$ enhances stability without backpropagating gradients, relying on the EMA strategy of $g_{\theta}$. During training, the source image $x_{S}^{(m)}$ and label $y_{S}^{(m)}$ train $g_{\theta}$, while the target image $x_{T}^{(n)}$ generates pseudo-labels $p_{T}^{(n)}$ through $t_{\theta}$, facilitating self-training. Data augmentation creates two views for contrastive learning, enriching training signals.
As illustrated in Figure 2, the network $g_{\theta}$ is a model specifically designed for image segmentation tasks. Its architecture comprises a feature extraction backbone $f$ and a decoding head responsible for segmentation predictions. To enhance the model’s stability and performance, we introduce an Exponential Moving Average (EMA) teacher network $t_{\theta}$, which has an identical architecture to $g_{\theta}$ but does not backpropagate gradients during training. This design ensures that the weight updates of $t_{\theta}$ rely solely on the exponential moving average strategy of $g_{\theta}$, effectively reducing fluctuations during the training process.
@@ III-F SiamSeg Network Architecture
During training, as indicated by the blue arrows in the figure, the source image $x_{S}^{(m)}$ and its corresponding source label $y_{S}^{(m)}$ are utilized to initially train the network $g_{\theta}$ through supervised learning. In the step denoted by the black arrows, the target image $x_{T}^{(n)}$ is processed by the EMA teacher network $t_{\theta}$, generating pseudo-labels $p_{T}^{(n)}$ to replace the inaccessible target labels $y_{T}^{(n)}$ during training. This process not only reduces the reliance on true labels but also enhances the model’s adaptability in the target domain. After generating pseudo-labels, the target image $x_{T}^{(n)}$ and pseudo-label $p_{T}^{(n)}$ are combined for Self-Training. It is important to emphasize that during the pseudo-label generation process, the teacher network $t_{\theta}$ does not participate in backpropagation due to the introduction of the stop-gradient operation.
Furthermore, by applying data augmentation to the target image $x_{T}^{(n)}$, we generate two distinct views, which together form positive sample pairs and provide rich training signals, with contrastive learning, for model. The following section III-F2 will explore the specific architecture and implementation details of the contrastive learning network, revealing its significant role in self-supervised learning.
III-F2 detail of contrastive learning
Figure 3: Siamese Network Architecture for Contrastive Learning. This figure illustrates the architecture of the Siamese network used for contrastive learning. The network consists of two identical sub-networks that share the same model weights, ensuring consistency in feature extraction. Each input image $x_{T}^{(n)}$ undergoes various data augmentation techniques (eg., Resize Crop, Color Jitter, Gray Scale, Gaussian Blur, Filp), generating two distinct views: $x_{T1}^{(n)}$ and $x_{T2}^{(n)}$. The first view is processed through an encoder and a prediction MLP head, producing output $\mathbf{p}_{1}^{(n)}$. In contrast, the second view is processed only through the encoder, yielding output $\mathbf{z}_{2}^{(n)}$ while employing a stop-gradient operation to optimize the network. The diagram highlights the symmetrical workflows of both views and the optimization process aimed at maximizing the similarity between $\mathbf{p}_{1}^{(n)}$ and $\mathbf{z}_{2}^{(n)}$, thereby enhancing feature representation for effective contrastive learning.
@@
The implementation of contrastive learning utilizes the Siamese network architecture, a simple yet effective strategy [39] [41] [44]. As shown in Fig. 3, the Siamese network comprises two identical sub-networks that share the same model weights, ensuring consistency during feature extraction and facilitating effective contrastive learning. Specifically, an input image $x_{T}^{(n)}$ undergoes data augmentation (eg., Resize Crop, Color Jitter, Gray Scale, Gaussian Blur, Filp), generating two distinct views, $x_{T1}^{(n)}$ and $x_{T2}^{(n)}$. These views are processed by the contrastive networks. Despite sharing parameters, their outputs differ slightly. In the upper workflow, the view $x_{T1}^{(n)}$ is first encoded by the encoder $E$ and then processed through the prediction MLP head $h$ to yield output $\mathbf{p}_{1}^{(n)}$. In contrast, the view $x_{T2}^{(n)}$ is also passed through the encoder to produce output $\mathbf{z}_{2}^{(n)}$ but does not go through the prediction MLP head; instead, it employs a stop-gradient operation. To maximize the similarity between outputs $\mathbf{p}_{1}^{(n)}$ and $\mathbf{z}_{2}^{(n)}$, we optimize using Equation 7. This process encourages the model to fully exploit the similarities between positive sample pairs, thereby enhancing the effectiveness and robustness of feature representation. The lower workflow mirrors the upper one, ensuring the unity and consistency of the overall contrastive learning process. Through this structured design, the Siamese network effectively performs feature learning, providing strong representational capabilities for our RS tasks.
@@ IV-A1 Potsdam and Vaihingen datasets
The Potsdam and Vaihingen datasets are part of the ISPRS 2D semantic segmentation benchmarks [50]. The Potsdam (POT) dataset comprises $38$ remote sensing images with a resolution of $6000\times 6000$ pixels and a ground sampling distance of $5$ meters. It features three different image modalities: IRRG, RGB, and RGBIR, where the first two modalities have three channels, and the last one has four. In this study, we primarily utilize the first two modalities. The Vaihingen (VAI) dataset contains $33$ remote sensing images with resolutions ranging from $1996\times 1996$ to $3816\times 2550$ pixels, with a ground sampling distance of $9$ centimeters. This dataset includes only one image modality (IRRG). Both datasets share six common classes: impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background, where the clutter/background class includes water boundaries and other irrelevant objects for semantic segmentation. In our study, we utilize the IRRG and RGB image modalities. Notably, the visual appearance of some objects may differ in the IRRG modality compared to conventional images; for instance, trees and low vegetation appear red.
To reduce computational overhead, we cropped the images to a smaller size of $512\times 512$ pixels. For the POT and VAI datasets, we used cropping strides of $512$ and $256$, resulting in $4598$ and $1696$ images, respectively. Subsequently, we split the POT and VAI datasets into training and testing sets, following previous wokrs [13] [51]. In the POT dataset, the number of images in the training and testing sets is $2904$ and $1694$, respectively, while in the VAI dataset, these numbers are $1296$ and $440$.
Due to the urban nature of Potsdam, it contains a higher number of vehicles, whereas Vaihingen includes more residential and agricultural areas. Although both datasets have the same task objectives and label space, they exhibit significant differences in feature distribution, making them well-suited for evaluating the performance of unsupervised domain adaptation (UDA) methods. Accordingly, we established four cross-domain remote sensing semantic segmentation tasks:
@@ IV-B Evaluation Metrics
In this study, we adopt the F1-score and Intersection over Union (IoU) as evaluation metrics, following previous methods in the field of remote sensing (RS) semantic segmentation domain adaptation (UDA). The specific calculation formulas are as follows:
@@
In equations (9) and (10), $TP$ represents true positives, $FP$ denotes false positives, and $FN$ indicates false negatives. The IoU, also known as the Jaccard index, and the F1-score, referred to as the Dice coefficient, effectively reflect the accuracy and reliability of model performance in semantic segmentation tasks. These metrics are particularly valuable when addressing imbalanced datasets, as the F1-score provides a more comprehensive evaluation of performance. By utilizing these evaluation metrics, we can systematically compare the performance of different methods in cross-domain RS image segmentation tasks.
@@ IV-C2 Network Architecture Details
Considering the outstanding performance of the Transformer-based Segformer [18] in semantic segmentation tasks, we chose Mix Vision Transformers (MiT) [18] as the backbone network (denoted as backbone $f$). This model was pre-trained on ImageNet [56], eliminating the need to use convolutional neural network (CNN)-based architectures like DeepLabV3+ [57]. In semantic segmentation, capturing both global and local features is critical, and feature fusion strategies are often used to improve segmentation performance. For this reason, we adopt the context-aware multi-scale feature fusion decoder head designed by Hoyer et al. [33], given its superior performance in UDA-based semantic segmentation. The feature pyramid generated by the backbone $f$ has channel sizes of [64, 128, 320, 512]. The decoder head projects all features to 256 channels while maintaining a resolution of $\frac{H}{4}\times\frac{W}{4}$ for further processing and segmentation tasks.
@@ IV-C3 Training
In the pseudo-label generation phase, the temperature parameter $\tau$ in Equation 2 is set to 0.999. The total number of training iterations is 40,000, with a batch size of 12, containing 6 source domain and 6 target domain images. During the preprocessing of the Potsdam (POT) and Vaihingen (VAI) datasets, image scaling is set to $img\_scale=(512,512)$, and the aspect ratio range is set to $ratio\_range=(1.0,2.0)$. Crops are sized at $512\times 512$, with $cat\_max\_ratio=0.75$ and a flipping probability of 0.5. Normalization is performed using the mean values [123.675, 116.28, 103.53] and standard deviations [58.395, 57.12, 57.375].
For the LoveDA dataset, we retain the settings from POT and VAI but adjust certain parameters. Specifically, image scaling is set to $img\_scale=(1024,1024)$ and the aspect ratio range to $ratio\_range=(0.5,1.5)$. During the contrastive learning process, data augmentations (Sim. Aug.) use a crop size of $size=(512,512)$ and scale range $scale=(0.6,1.0)$. Color jitter parameters are set to a brightness, contrast, saturation, and hue of 0.25 each, with a random application probability of 0.6. Grayscale has a probability of 0.2, Gaussian blur has a probability of 0.5, and both horizontal and vertical flips are applied with a probability of 0.5. These settings aim to enhance model generalization through data augmentation, improving performance in cross-domain remote sensing image segmentation tasks.
IV-C4 Testing and Validation
For the testing and validation stages, we use MultiScaleFlipAug [54] to evaluate semantic segmentation performance, applying multi-scale augmentation and flipping during testing. For the POT and VAI datasets, since the image size is set to $512\times 512$ after dataset preparation, we set $img\_scale=(512,512)$ and use an image scaling ratio of $img\_ratios=[1.0]$. For the LoveDA dataset, where the original image size is $1024\times 1024$, the image scaling is set to $img\_scale=(1024,1024)$ with scaling ratios $img\_ratios=[0.75,1.0,1.25,1.5]$.
All experiments are conducted within the MMSegmentation framework to ensure consistency and reproducibility. Moreover, all model training is performed using Nvidia A100 GPU×$\times 4$, enhancing training efficiency and performance. These settings provide a solid foundation for the subsequent experimental results, ensuring the validity and stability of the model across different datasets.
@@ IV-D1 Quantitative Results
POT IRRG → VAI IRRG. In this task, Potsdam IRRG is used as the source domain, while Vaihingen IRRG is the target domain. POT contains 2904 labeled training images, and VAI contains 1296 unlabeled training images. Additionally, 440 labeled test images from VAI are used to evaluate model performance. The results are shown in Table I.
Even when trained only on the source domain, Segformer outperforms several UDA methods based on the DeeplabV3+ [57] series, such as AdaptSegNet, ProDA, and DualGAN, in certain categories. The SimSeg method, proposed in this paper, achieves a nearly 20% improvement in mIoU and a 17% increase in mF1, thanks to the combination of pseudo-labeling self-training (ST) and contrastive learning (C.L.).
SimSeg, through the use of pseudo-label ST for UDA, surpasses the performance of adversarial learning methods like CIA-UDA and CNT, highlighting ST as a highly competitive approach. Moreover, compared to SiamSeg without C.L., SiamSeg with C.L. exhibits a further 2.53% and 2.61% improvement in mIoU and mF1, respectively. This demonstrates the effectiveness of contrastive learning in providing additional supervision signals, thereby enabling the backbone network to better learn target domain features.
@@
POT RGB → VAI IRRG. In this task, we utilized Potsdam RGB as the source domain and Vaihingen IRRG as the target domain. The POT dataset contains 2,904 labeled training images, while the VAI dataset has 1,296 unlabeled training images and 440 labeled testing images for performance evaluation. The results, shown in Table II, exhibit a trend similar to the POT IRRG → VAI IRRG task.
Notably, Segformer, as the baseline method, is trained solely on the source domain (POT RGB) and tested directly on the target domain (VAI IRRG). Although Segformer shows relatively balanced performance across all categories, its generalization capability in cross-domain scenarios is limited, achieving an mIoU of 36.95 and an mF1 of 48.17, indicating performance bottlenecks in domain transfer.
For the adversarial learning-based UDA methods, AdaptSegNet and ProDA display significant shortcomings across multiple categories, particularly with AdaptSegNet performing poorly in the “car” and “low vegetation” categories. Although ProDA shows slight improvements in these categories, its overall mIoU and mF1 scores remain unsatisfactory, suggesting inadequate adaptability for remote sensing image semantic segmentation tasks.
Among UDA methods designed for the remote sensing domain, DualGAN demonstrates noteworthy performance in the “car” and “low vegetation” categories, while DNT showcases strong generalization capability, especially in the “building” and “impervious surface” categories. CIA-UDA exhibits excellent performance in the “car” and “tree” categories and displays robust cross-domain adaptability across several classes.
The SiamSeg method without contrastive learning performs well in the “building” category. However, the version with contrastive learning exhibits significant improvements across multiple categories, particularly in “low vegetation” and “impervious surface,” confirming that contrastive learning provides additional supervisory signals that enhance model performance on the target domain.
TABLE II: Cross-domain RS image semantic segmentation results from Potsdam RGB to Vaihingen IRRG. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
@@
VAI IRRG → POT IRRG. In this task, we used Vaihingen IRRG as the source domain and Potsdam IRRG as the target domain. The VAI dataset contains 1,296 labeled training images, while the POT dataset comprises 2,904 unlabeled training images, along with 1,694 labeled POT testing images used for performance evaluation of each model. The experimental results are presented in Table III.
In this task, Segformer serves as the baseline method, trained solely on the source domain VAI IRRG and tested on the target domain POT IRRG. Despite its strong performance in the ”building” category, the overall results remain weak, indicating a significant drop in performance during domain transfer.
AdaptSegNet performs inadequately in the ”car” category, while ProDA shows slight improvement in this category and achieves higher IoU in the ”tree” and ”impervious surface” categories. However, both adversarial learning methods maintain a low to moderate overall mIoU, suggesting limited cross-domain adaptability for this task.
DualGAN exhibits some improvement in the ”car” and ”tree” categories compared to AdaptSegNet and ProDA, yet it lacks competitive generalization capabilities and performs averagely in cross-domain adaptation tasks. In contrast, CIA-UDA excels across multiple categories, particularly in ”car” and ”impervious surface,” demonstrating significant advantages. DNT exhibits stable performance across most categories, showcasing reliable cross-domain generalization capabilities.
SiamSeg, by incorporating contrastive learning, significantly enhances performance across all categories, particularly excelling in ”car,” ”building,” and ”impervious surface.” Its mIoU and mF1 scores are the highest among all methods, confirming the effectiveness of contrastive learning in cross-domain semantic segmentation tasks.
TABLE III: Cross-domain RS image semantic segmentation results from Vaihingen IRRG to Potsdam IRRG. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
@@
VAI IRRG → POT RGB. In this task, we selected Vaihingen IRRG as the source domain and Potsdam RGB as the target domain. The Vaihingen dataset (VAI) contains a total of 1,296 labeled training images, while the Potsdam dataset (POT) includes 2,904 unlabeled training images. Additionally, the POT dataset contains 1,694 labeled testing images used to evaluate the cross-domain performance of each model. The experimental results are detailed in Table IV.
Overall, SiamSeg demonstrates superior performance in cross-domain adaptation tasks, particularly excelling in the ”car,” ”tree,” and ”building” categories, achieving the best IoU and F1 scores. The incorporation of contrastive learning effectively enhances the model’s generalization capability in the target domain, significantly improving cross-domain segmentation performance. In contrast, CIA-UDA also shows strong competitiveness, particularly in the ”car” and ”building” categories, achieving results close to those of SiamSeg, indicating its excellent transferability in certain categories.
Conversely, other adversarial learning-based methods, such as ProDA and DualGAN, display relatively weaker overall performance. Although they show improvements in some categories, they fail to consistently enhance the cross-domain segmentation capability across all categories. For example, ProDA performs well in the ”low vegetation” category but demonstrates limited performance in the ”building” and ”impervious surface” categories. Similarly, DualGAN, while achieving robust IoU in the ”tree” and ”building” categories, still underperforms in others, lacking consistent cross-domain adaptability.
The results in Table IV further validate the cross-domain advantages of SiamSeg following the introduction of contrastive learning, particularly in multi-class segmentation tasks, where mIoU and mF1 metrics are significantly improved. Meanwhile, CIA-UDA, as a competitor, also exhibits good cross-domain generalization capabilities, especially in key target categories. In comparison, traditional adversarial learning methods show uneven performance, with notable limitations in cross-domain segmentation effectiveness.
TABLE IV: Cross-domain RS image semantic segmentation results from Vaihingen IRRG to Potsdam RGB. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
@@
In this task, Rural and Urban serve as the source and target domains, respectively. The Rural domain contains 1,296 labeled training images, while the Urban domain includes 1,366 unlabeled training images. Additionally, 677 labeled testing images from the Urban domain were used to evaluate the performance of each model. The experimental results are presented in Table V.
From the results, AdaptSegNet shows good performance in the ”building” category, but overall, it performs weakly with an mIoU of only 32.68. Its performance is particularly poor in complex scenes, such as the ”barren” class, with an IoU of only 13.62, indicating limited generalization capability. FADA achieves an overall mIoU of 31.41, highlighting its limitations in processing unlabeled target domain data, particularly in segmentation tasks involving complex terrains and vegetation. The CLAN method exhibits relatively stable performance across multiple categories with an mIoU of 33.11; however, it underperforms in the ”Forest” category, suggesting that its generalization in natural scenes requires improvement.
PyCDA demonstrates strong cross-domain segmentation capabilities with an mIoU of 36.25, excelling particularly in the ”Forest” and ”Road” categories, showcasing adaptability in diverse scenarios. CBST stands out with an mIoU of 41.32, indicating excellent performance across multiple categories, especially in ”Road” and ”Building.” IAST achieves an IoU of 86.01 in the ”Water” category, the highest among all comparative methods, resulting in an overall mIoU of 40.48 and demonstrating strong performance across various categories. DCA achieves an mIoU of 46.36, reflecting its robust cross-domain segmentation ability.
The proposed SiamSeg method outperforms all other methods with an impressive mIoU of 51.72, particularly excelling in the ”Building,” ”Road,” and ”Forest” categories, showcasing exceptional cross-domain generalization capability. Compared to other methods, SiamSeg integrates contrastive learning into self-supervised training, offering a novel approach for unsupervised domain adaptation (UDA) in semantic segmentation, and significantly enhancing the model’s performance in the target domain, achieving highly competitive results.
TABLE V: ross-domain RS image semantic segmentation results from LoveDA rural to urban. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.

### v3

@@ III-C Contrastive Learning
$L_{CLR}^{(i)}=\frac{1}{2}\cdot D(p_{1}^{(i)},sg(z_{2}^{(i)}))+\frac{1}{2}\cdot D(p_{2}^{(i)},sg(z_{1}^{(i)})).$ (7)
In the loss function, the stop-gradient operation is a crucial step, treating the variables in $sg(\cdot)$ as constants. This operation effectively prevents representation collapse [46]. Specifically, in the first term of Equation (7), the encoder processing $x_{T_{2}}^{(i)}$ does not receive gradients from $z_{2}^{(i)}$, while in the second term, the encoder processing $x_{T_{1}}^{(i)}$ does receive gradients from $p_{2}^{(i)}$. This design effectively enhances the model’s representation learning capabilities, improving its performance in the target domain.
III-D Proposed UDA Losses
In a given training step, we acquire the $m$-th image and its corresponding label from the source domain, represented as $x_{S}^{(m)}$ and $y_{S}^{(m)}$. Simultaneously, we obtain the $m$-th image from the target domain and its pseudo-label generated by the teacher network $t_{\theta}$, denoted as $x_{T}^{(n)}$ and $p_{T}^{(n)}$. The total loss function we mathematically define is as follows:
$L_{total}^{(m;n)}=L_{S}^{(m)}+\beta\cdot L_{T}^{(n)}+\gamma\cdot L_{CLR}^{(n)}.$ (8)
@@ III-D Proposed UDA Losses
Specifically, the loss $L_{S}^{(m)}$ guides the model’s performance on source domain data, while $L_{T}^{(n)}$ leverages pseudo-label to drive learning in the target domain, further enhancing the model’s generalization ability in this domain. Finally, $L_{CLR}^{(n)}$ introduces richer feature representations through contrastive learning, enabling the model to better adapt to potential differences between the source and target domains. By integrating these losses, we can effectively reduce the domain gap between the source and target domains, thereby improving the model’s performance on the unlabeled target domain data.
III-E SiamSeg Network Architecture
III-E1 overall
Figure 2: Overall of SiamSeg. The network $g_{\theta}$ is designed for image segmentation and comprises a feature extraction backbone $f$ and a decoding head, an EMA teacher network $t_{\theta}$ and a contrastive network.
Figure 3: Detail of Contrastive Network. This figure illustrates the architecture of the Siamese network used for contrastive learning. The network consists of two identical sub-networks that share the same model weights, ensuring consistency in feature extraction.
As illustrated in Figure 3, the network $g_{\theta}$ is a model designed for image segmentation tasks. Its architecture comprises a feature extraction backbone $f$ and a decoding head responsible for segmentation predictions. To enhance the model’s stability and performance, we introduce an EMA teacher network $t_{\theta}$, which has an identical architecture to $g_{\theta}$ but does not backpropagate gradients during training.
@@ III-E SiamSeg Network Architecture
During training, as indicated by the blue arrows in the figure, the source image $x_{S}^{(m)}$ and its corresponding source label $y_{S}^{(m)}$ are utilized to initially train the network $g_{\theta}$ through supervised learning. In the step denoted by the black arrows, the target image $x_{T}^{(n)}$ is processed by the EMA teacher network $t_{\theta}$, generating pseudo-label $p_{T}^{(n)}$ to replace the inaccessible target labels $y_{T}^{(n)}$ during training. Furthermore, by applying data augmentation to the target image $x_{T}^{(n)}$, we generate two distinct views, which together form positive sample pairs and provide rich training signals, with contrastive learning, for model.
III-E2 detail of contrastive learning
@@
The implementation of contrastive learning utilizes the Siamese network architecture, a simple yet effective strategy [41, 43, 46]. As shown in Fig. 3, the contrastive learning network comprises two identical sub-networks that share the same model weights, ensuring consistency during feature extraction and facilitating effective contrastive learning. Specifically, an input image $x_{T}^{(n)}$ undergoes data augmentation (eg., Resize Crop, Color Jitter, Gray Scale, Gaussian Blur, Filp), generating two distinct views, $x_{T1}^{(n)}$ and $x_{T2}^{(n)}$. These views are processed by the contrastive networks. Despite sharing parameters, their outputs differ slightly. In the upper workflow, the view $x_{T1}^{(n)}$ is first encoded by the encoder $E$ and then processed through the prediction MLP head $h$ to yield output $\mathbf{p}_{1}^{(n)}$. In contrast, the view $x_{T2}^{(n)}$ is also passed through the encoder to produce output $\mathbf{z}_{2}^{(n)}$ but does not go through the prediction MLP head; instead, it employs a stop-gradient operation. To maximize the similarity between outputs $\mathbf{p}_{1}^{(n)}$ and $\mathbf{z}_{2}^{(n)}$, we optimize using Equation 7. This process encourages the model to fully exploit the similarities between positive sample pairs, thereby enhancing the effectiveness and robustness of feature representation. The lower workflow mirrors the upper one, ensuring the unity and consistency of the overall contrastive learning process. Through this structured design, the Siamese network effectively performs feature learning, providing strong representational capabilities for our RS tasks.
@@ IV-A1 Potsdam and Vaihingen datasets
The Potsdam and Vaihingen datasets are part of the ISPRS 2D semantic segmentation benchmarks [48]. The Potsdam (POT) dataset comprises $38$ remote sensing images with a resolution of $6000\times 6000$ pixels and a ground sampling distance of $5$ meters. It features three different image modalities: IRRG, RGB, and RGBIR, where the first two modalities have three channels, and the last one has four. In this study, we primarily utilize the first two modalities. The Vaihingen (VAI) dataset contains $33$ remote sensing images with resolutions ranging from $1996\times 1996$ to $3816\times 2550$ pixels, with a ground sampling distance of $9$ centimeters. This dataset includes only one image modality (IRRG). Both datasets share six common classes: impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background.
To reduce computational overhead, we cropped the images to a smaller size of $512\times 512$ pixels. For the POT and VAI datasets, we used cropping strides of $512$ and $256$, resulting in $4598$ and $1696$ images, respectively. Subsequently, we split the POT and VAI datasets into training and testing sets, following previous wokrs [14, 49]. In the POT dataset, the number of images in the training and testing sets is $2904$ and $1694$, respectively, while in the VAI dataset, these numbers are $1296$ and $440$. We established four cross-domain remote sensing semantic segmentation tasks:
@@ IV-B Evaluation Metrics
In this study, we adopt the F1-score (F1) and Intersection over Union (IoU) as evaluation metrics, following previous methods in the field of remote sensing (RS) semantic segmentation domain adaptation (UDA). The specific calculation formulas are as follows:
@@
In equations (9) and (10), $TP$ represents true positives, $FP$ denotes false positives, and $FN$ indicates false negatives. The IoU, also known as the Jaccard index, and the F1-score, referred to as the Dice coefficient, effectively reflect the accuracy and reliability of model performance in semantic segmentation tasks.
@@ IV-C2 Network Architecture Details
Considering the outstanding performance of the transformer-based Segformer [19] in semantic segmentation tasks, we chose Mix Vision Transformers (MiT) [19] as the backbone network $f$. This model was pre-trained on ImageNet [54]. In semantic segmentation, capturing both global and local features is critical, and feature fusion strategies are often used to improve segmentation performance. For this reason, we adopt the context-aware multi-scale feature fusion decoder head designed by Hoyer et al. [35], given its superior performance in UDA-based semantic segmentation.
@@ IV-C3 Training
In the pseudo-label generation phase, the temperature parameter $\tau$ in Equation 2 is set to 0.999. The total number of training iterations is 40,000, with a batch size of 12, containing 6 source domain and 6 target domain images.
During the contrastive learning process, data augmentations (Sim. Aug.) use a crop size of $size=(512,512)$ and scale range $scale=(0.6,1.0)$. Color jitter parameters are set to a brightness, contrast, saturation, and hue of 0.25 each, with a random application probability of 0.6. Grayscale has a probability of 0.2, Gaussian blur has a probability of 0.5, and both horizontal and vertical flips are applied with a probability of 0.5. These settings aim to enhance model generalization through data augmentation, improving performance in cross-domain remote sensing image segmentation tasks.
All experiments are conducted within the MMSegmentation framework to ensure consistency and reproducibility. Moreover, all model training is performed using Nvidia A100 GPU×$\times 4$, enhancing training efficiency and performance.
@@ IV-D1 Quantitative Results
From Table I, II, III and IV, we can find that for some of the methods that use Deeplab as a backbone, such as AdaptSegNet, ProDA and DualGAN, the IoU is lower than the prediction accuracies obtained using Segformer, which is trained only on the source domain, in most cases. This is due to the fact that Segformer is a transformer-based method, while DeepLab is a convolution-based method. For the more complex image features of RS images, the attention mechanism of transformer is able to better utilize the feature context of the image, so some of the methods still have lower accuracy when facing RS images, even for UDA methods. As for CIA-UDA and DNT, which are methods designed for the characteristics of RS images, they achieve competitive performance in segmentation results. For the methods Siamseg and SiamSeg without C.L. in this paper, without the addition of contrastive learning, the reliance on ST still achieves good performance, which shows that the ST method is an effective UDA method. However, since the domain shift of RS images varies greatly between different domains, this leads to insufficient learning of the target domain image. And after adding contrastive learning, it can be observed that the method predicts a significant increase in IoU. Thus indicating that the addition of contrastive learning does allow the model to learn more image features of the target domain, thus greatly enhancing the performance.
@@
TABLE II: Cross-domain RS image semantic segmentation results from Potsdam RGB to Vaihingen IRRG.
@@
TABLE III: Cross-domain RS image semantic segmentation results from Vaihingen IRRG to Potsdam IRRG.
@@
TABLE IV: Cross-domain RS image semantic segmentation results from Vaihingen IRRG to Potsdam RGB.
@@
The experimental results are presented in Table V. LoveDA dataset has more number of classifications and more differences between rural and urban,so most of the methods have low IoU on this dataset. Adversarial learning based methods, such as AdaptSegNet and CLAN, all perform very poorly because they cannot reduce the domain gap well. While methods using ST, such as CBST, IAST and SiamSeg w./o. C.L. have good performance, which suggests that ST is a good choice for solving cross-domain problems. However, when facing the more complex cross-domain remote sensing images, the ST method is not able to fully learn the features of the target domain due to the large domain gap, while the siamseg with the introduction of contrastive learning learns the features of the target domain better thanks to the contrastive learning, and thus enhances the performance.
TABLE V: ross-domain RS image semantic segmentation results from LoveDA rural to urban.